solve the following system by addition. if a unique solution does not exist, state whethere the system is consistent or dependent.
2x+3y=1
5x+3y=16

Answer by elima(1433)

2x+3y=1
5x+3y=16
=============
We will start by multiplying the second equation with -1;
(-1)5x+3y=16
-5x-3y=-16
===================
Now add the equations together;
2x+3y=1
-5x-3y=-16
----------
-3x=-15
x=5
=================
Now lets plug 5 into one of the equations for x;
2(5)+3y=1
10+3y=1
3y=-9
y=-3
:)